OpenStudy (anonymous):

An identification code is to consist of five letters followed by three digits. How many different codes are possible if repetition is permitted?

OpenStudy (anonymous):

i am not very good at these type of problems.. not sure how to do them

OpenStudy (anonymous):

For the first letter you have 24 different choices. The same is true for all the letters and you have 10 choices for the number: 24x24x24x24x24x10x10x10 should give you the result.

OpenStudy (anonymous):

where did u get the 24?

OpenStudy (anonymous):

oops, I meant 26 (letters in the alphabet) :P. 26x26x26x26x26x10x10x10

OpenStudy (anonymous):

130,000 ways?

OpenStudy (anonymous):

It should be 1.1881376 x 10^10

OpenStudy (anonymous):

\[26^5 \cdot 10^3 = 11,882,376\]

OpenStudy (anonymous):

hmmm i did my math totally wrong.. lol

OpenStudy (anonymous):

26^5 = 11,881,376 but then you have to multiply by 10^3

OpenStudy (anonymous):

oh ok.. thank you :)